OpenStudy (anonymous):

Find the derivative of the function. f(t) = t^3/2 log9(sqrt(t + 5)) the 9 is a small srcipt after the log.

OpenStudy (zarkon):

this \[f(t)=\frac{t^3}{2\log_9(\sqrt{t+5})}\]

OpenStudy (anonymous):

no its t^2/3

OpenStudy (anonymous):

3/2*

OpenStudy (anonymous):

sorry its t^3/2 all over the log stuff

OpenStudy (zarkon):

\[f(t)=t^{3/2}\log_9(\sqrt{t+5})\]

OpenStudy (anonymous):

yeh

OpenStudy (zarkon):

\[=f(t)=t^{3/2}\frac{1}{2}\log_9(t+5)\] then \[f'(t)=\frac{3}{4}t^{1/2}\log_9(t+5)+\frac{1}{2}t^{3/2}\frac{1}{t+5}\frac{1}{\ln(9)}\]

OpenStudy (anonymous):

\[\frac{t^{3/2}}{\log_{9}\sqrt{t+5}} \]?

OpenStudy (anonymous):

what he got looks good

OpenStudy (anonymous):

no its t^(2/3)

OpenStudy (anonymous):

so change the t^3/2 to a 2/3 and do the product rule

OpenStudy (zarkon):

\[f(t)=t^{2/3}\frac{1}{2}\log_9(t+5)\] :)

OpenStudy (anonymous):

why si the derivtive so long..somehtign looks wrong:S

OpenStudy (anonymous):

1/2 log .. you sure that means logbase9(t+5)/2

OpenStudy (anonymous):

whatever zarkon posted thats exactlly how my problem looks

OpenStudy (anonymous):

why don't you take a picture of it

OpenStudy (anonymous):

zarkon?..whatever zarkon posted is idetical to my problem

OpenStudy (anonymous):

so it has a 1/2 in front/

OpenStudy (zarkon):

ok...then just modify my first answer

OpenStudy (anonymous):

all he's simply doing is the product rule

OpenStudy (anonymous):

these are the rules

OpenStudy (anonymous):

ll here is the rule for the last part the first part is simple and you are just going to use the product rule whichis d/dx[uv] = uv'+vu'

OpenStudy (anonymous):

i'm not sure how he got u' on the bottom tho

OpenStudy (zarkon):

where?

OpenStudy (anonymous):

above for the derivative of logbase9(t+5)

OpenStudy (anonymous):

o wait u' = 1 nvm

OpenStudy (anonymous):

zarkon can u please give me a finla result im so confused...i only have 7 mins left:'(

OpenStudy (anonymous):

the answer is above it'd be 2/3 * 1/6

OpenStudy (anonymous):

instead

OpenStudy (anonymous):

and instead of t^1/2 it'll be t^-1/2

OpenStudy (anonymous):

can u write out a final asnwer for me please..my comp keeps freezing

OpenStudy (anonymous):

2/3*1/2 which would be 1/3 i meant above

OpenStudy (zarkon):

\[t^{-1/3}\]

OpenStudy (anonymous):

yes i'm just fumbling aroudn becaause i'm using the numbers from the other equations >_<

OpenStudy (anonymous):

zarkonn can u please give me a finla answer..like type it out..my chat screen looks all weird:(

OpenStudy (zarkon):

\[f(t)=t^{2/3}\frac{1}{2}\log_9(t+5)\] \[f'(t)=\frac{1}{2}\left[\frac{2}{3}t^{-1/3}\log_9(t+5)+t^{2/3}\frac{1}{t+5}\frac{1}{\ln(9)}\right]\]

OpenStudy (anonymous):

what he has is right

OpenStudy (zarkon):

I know ;)

OpenStudy (anonymous):

hahaha but you could simplify the back

OpenStudy (zarkon):

na...I always tell my students to never simplify.

OpenStudy (zarkon):

:)

OpenStudy (anonymous):

you're a teacher

OpenStudy (anonymous):

?

OpenStudy (zarkon):

professor

OpenStudy (anonymous):

at?

OpenStudy (zarkon):

some college/university in the USA :)